OpenStudy (anonymous):

***URGENT HELP*** ***Medal and Fan Reward*** Light traveling through air at 3.00 · 10^8 m/s reaches an unknown medium and slows down to 2.00 · 10^8 m/s. What is the index of refraction of that medium? n =

OpenStudy (anonymous):

When ight traveling through air at 3.00 · 108 m/sec. reaches an unknown medium and slows down to 2.00 · 108 m/sec. The index of refraction is 1.

OpenStudy (anonymous):

The index of refraction, n, is a ratio that compares the speed of light in a vacuum to the speed of light in another medium. n = c/v where c = speed of light in a vacuum = 3 X 108 m/s The speed of light in all other media is less than in a vacuum. Therefore, n is always larger than one. It may be thought of as the measure of the slowness of light in a medium: the slower light travels in medium, the larger n will be for that medium. Since n divides speed by speed it has no units. The frequency of light does not change as it enters a different medium, but its wavelength does. Since v = fL where f = frequency and L = wavelength then v1/v2 = L1/L2

OpenStudy (nathan917):

Copy and pasted that...

OpenStudy (anonymous):

should I get 0.667?

OpenStudy (anonymous):

Its an example. Not hard to help her by giving an example to help work out the problem

OpenStudy (anonymous):

or 1.5?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

The approximate speed of light in a vacuum is found to be 3.00 x 108 m/s. The speed of light is designated as c rather than v and is related to wavelength and frequency by the following equation, c = f8 A light source, or body that emits light, is termed luminous. A body that reflects light is illuminated. The rate at which a body emits light depends on the rate at which it puts energy into it. This rate is called luminous flux (P) and the unit of luminous flux is the lumen (lm). The rate of light that falls on a unit surface area is called illuminance (E). The relationship between the flux and the illuminance is E = P/4Bd2 where E is the illuminance, P represents the luminous flux, and d is the distance from the surface.

OpenStudy (anonymous):

I'm saying it should be 1.5 because I got 1

OpenStudy (anonymous):

it was 1.5.... thanks for the help :)

OpenStudy (anonymous):

Your welcome (: